OpenStudy (anonymous):

find the distance between (-1,7) (-4,3)

OpenStudy (anonymous):

use distance formual: \[\sqrt{(x1-x2)^2+(y1-y2)^2}\]

OpenStudy (anonymous):

5

OpenStudy (anonymous):

c^2= a^2+b^2 c=sqr-root[(3-7)^2 + (-4-(-1))^2] c=sqr-root{25] c=5 therefore the distance is 5.